# Towards Scalability and Extensibility of Query Reformulation Modeling in E-commerce Search

### v1

@@ 4.3. Model training with ANCE mining
Previous research has introduced the concept of Approximate Nearest Neighbor Negative Contrastive Estimation (ANCE) (Sohn, 2016) to address the inefficacy of using in-batch negative in contrastive learning. ANCE entails the incorporation of hard negatives, which are selected from the entire dataset for each query within a batch. This selection process is carried out through a sequence of retrieval model inferences and KNN indexing. In the context of QR, we define the set of hard negatives, denoted as $H(q_{i})$, as the queries that are retrieved for $q_{i}$ but do not share any co-purchases with $q_{i}$.
We opt to utilize $B(q_{i})$ instead of $B_{filtered}(q_{i})$ when constructing the co-purchase dataset for ANCE. This choice is made to minimize the occurrence of strongly relevant pairs with sparse signal being inadvertently misused as hard negatives. We put forth distinct strategies to harness ANCE in the training of retrieval and re-ranking models. For the retrieval model, we employ a self-learning strategy, wherein the bi-encoder is iteratively finetuned with the ANCE dataset it generates. Conversely, for the re-ranking model, we adopt a learning-from-teacher strategy. Under this approach, the cross-encoder is finetuned using the ANCE dataset generated by the bi-encoder model. The ANCE mining and model continuously finetune process is described in Figure 3.
Figure 3. The ANCE mining and model continuous finetune process. Specifically, the filtered dataset is used for initial training of retrieval (bi-encoder) and reranking (cross-encoder) models. Then the bi-encoder model is iteratively finetuned using ANCE data derived by self-retrieved pairs subtracting pairs with co-purchase. At last, the cross-encoder model is finetuned with the ANCE data from the bi-encoder.
4.4. Refined training objectives
We use $s_{BE}(q_{i},q_{j})$ to represent the cosine similarity between L2 normalized embedding of $q_{i}$ and $q_{j}$, and $s_{CE}(q_{i},q_{j})$ to represent the similarity score given by a cross-encoder model. For retrieval task, the general training objective under the contrastive learning with in-batch negative framework is defined as
(3) $\mathcal{L}_{\mathrm{retrieval}}=-1/N\cdot\Sigma_{i,k}\frac{\exp{s_{BE}(q_{i},q_{ik})/\tau}}{\Sigma_{j\in\mathrm{batch}}\exp{s_{BE}(q_{i},q_{j})/\tau}}$
where $N$ is the number of positive query pairs, $\tau$ is a temperature parameter.
The combination of the contrastive learning framework and the infoNCE (Oord et al., 2018) (NT-XENT (Sohn, 2016)) loss has demonstrated superior effectiveness compared to point-wise training using loss such as mean-squared error (MSE) and cross-entropy (CE). However, this objective still exhibit a limitation in contrast to MSE or CE, as it treats every query pair uniformly and does not leverage the similarity score. While this characteristic does not hinder its performance in classic unsupervised contrastive learning tasks, where the positives generated through data augmentation are robust, it poses challenges in the context of QR, where the positives derived from data mining are prone to being false positives. Assigning equal importance to each pair can lead to suboptimal results. In this paper, we refine the loss formulation by adding a sample importance term as
(4) $\mathcal{L}_{\mathrm{retrieval,hard}}=-1/N\cdot\Sigma_{i,k}I_{JS}(q_{i},q_{ik})\cdot\\ \frac{\exp{s_{BE}(q_{i},q_{ik})/\tau}}{\Sigma_{j\in\mathrm{batch}}\exp{s_{BE}(q_{i},q_{j})/\tau}+\Sigma_{j\in\mathrm{sample}(H(q_{i}))}\exp{s_{BE}(q_{i},q_{j})/\tau}}$
where $I_{JS}(q_{i},q_{ik})$ equals to $1-JSD(dist(q_{i})||dist(q_{ik}))$. By doing so, the robustness of the training process is bolstered across varying levels of available resources. On one hand, in the presence of abundant training data, the impact of less pertinent pairs is mitigated, allowing for more focused learning. On the other hand, in scenarios with limited supervised resources, pairs that may not be strictly relevant-yet still bear some relevance-can still make a valuable contribution to the model’s training.
For re-ranking task, we use the cross-entropy loss
(5) $\mathcal{L}_{\mathrm{re-ranking}}=-I_{KL}(q_{i},q_{j})\cdot\log s_{CE}(q_{i},q_{j})-\\ (1-I_{KL}(q_{i},q_{j}))\cdot\log(1-s_{CE}(q_{i},q_{j}))$
where $I_{KL}(q_{i},q_{j})$ equals to $-KLD(dist(q_{j})||(dist(q_{i})+dist(q_{j}))/2$.
@@ 4.4. Refined training objectives
For cross-encoder training with hard negative, we group the pairs associated to each $q_{i}$ to form a positive set and employ pairwise loss rather than pointwise loss. The necessity for training in a pairwise manner arises particularly in the case of hard negatives, where positive and negative samples often exhibit only nuanced differences in their text representations. Pairwise training can enable the model to capture the supervised signal with greater precision, reduce variance during training, and thus, enhance the overall training efficacy. Specifically, we use circle loss (Sun et al., 2020), which is designed to direct the model’s attention toward ensuring that the logit for positive pairs surpasses that of negative pairs, rather than exclusively concentrating on the similarity score.
@@ 5.1. Dataset generation
We construct training and testing dataset with one-year anonymized Amazon search data. We first extract query-product pairs from anonymized historical customer behavior data. Then we perform self-join of these query-product pairs on product, and aggregate the joined result by product to obtain the query co-purchase data $D$. At last, we select 5,049 unique queries from $D$ and extract all query pairs containing the selected queries as testing data. The remaining pairs are divided into training and validation data, using a 90%/10% ratio. The statistics of the generated dataset is shown in Appendix A.
5.2. Backbone models
In order to boost training efficacy and efficiency, we developed both bi-encoder and cross-encoder models upon pre-trained language models (PLMs). These models employed a BPE tokenizer with a 300K vocabulary, trained using Amazon’s internal product data via the sentencepiece package (Kudo and Richardson, 2018).
@@ 5.3. Offline evaluation criteria
We employ distinct evaluation metrics for the retrieval and re-ranking tasks. For retrieval, we calculate the recall@100 and recall@1000 for both the top-3 paired queries and all queries paired to the given query in the testing data. This separation into top-3 and all queries is motivated by their significance in reflecting the model’s performance in downstream ranking and matching applications, respectively; while evaluating both recall@100 and recall@1000 aims to provide intuitions for determining the number of retrieved candidates. The paired queries are ranked based on the behavior-based similarity score, as described in Section 4.1. For queries with more than 100 or 1000 paired queries, we have slightly adjusted the recall formulation as follows
(7) $\frac{|q_{knn}\cap q_{paired}|}{|q_{paired}|}\rightarrow\frac{|q_{knn}\cap q_{paired}|}{min(|q_{paired}|,|q_{knn}|)}$
where $q_{knn}$ and $q_{paired}$ denote the set of retrieved queries and paired queries of q. Besides, we calculate both micro-recall, which represents the average recall of each query in the testing dataset, and macro-recall, which is derived by dividing the total number of retrieved paired queries by the total number of all paired queries.
@@ 5.4. Offline experiment result
We train a series of bi-encoder model and cross-encoder model, separately for retrieval and re-ranking tasks. The proposed approaches are progressively added upon a baseline model to study the individual benefit brought by each approach. The datasets used for training were constructed through query normalization-based augmentation. Additionally, we tested the models implemented under the current QR solution in Amazon, referred as ”baseline_previous”, whose only difference from baseline was that it employed the criteria defined by Equation 2 for mining positive labels.
Retrieval - The experimental results are detailed in Table 3. The baseline model (index rt2) was trained using contrastive loss (Equation 3). For simplicity, we will analyze the findings specifically for recall@100 in relation to the top-3 queries (in bold font), as the results for all metrics follow a similar trend. Notably, the inclusion of the sample importance term elevated recall@100 by 5.4% (index rt3 vs. rt2). This improvement was further augmented by the addition of ANCE, resulting in an incremental gain of 7.9% (index rt4 vs. rt3). Additionally, through successive iterations of the ANCE dataset and continuous model fine-tuning, recall@100 increased to 0.777 in round 2 (index rt5) and 0.798 in round 3 (index rt6). Note that we only perform three rounds of ANCE as we observed marginal benefit with more rounds. The outcomes suggest that the combined utilization of the sample importance term in the loss function and ANCE substantially enhances the retrieval model’s performance.
@@
Index Model Recall (top-3) Recall
@100 @1000 @100 @1000
micro macro micro macro micro macro micro macro
rt1 Baseline_previous 0.40798 0.41115 0.58694 0.59376 0.2747 0.20151 0.44657 0.34673
rt2 Baseline 0.59029 0.59322 0.92091 0.9304 0.47014 0.34508 0.87407 0.74717
rt3 +sample importance 0.64414 0.65069 0.92369 0.93217 0.51185 0.37882 0.86913 0.74003
rt4 +ANCE (round 1) 0.72323 0.72559 0.91881 0.9231 0.6075 0.46382 0.87634 0.74752
rt5 +ANCE (round 2) 0.77664 0.77706 0.93883 0.94461 0.65705 0.48759 0.89501 0.76242
rt6 +ANCE (round 3) 0.79837 0.79949 0.94361 0.94968 0.68076 0.51591 0.9038 0.77851
Re-ranking - the model performance is presented in Table 4. The baseline model was trained using cross-entropy loss. Surprisingly, our findings revealed a rather unexpected outcome: as the model’s focus shifted towards hard negative samples through the incorporation of ANCE and the adoption of circle loss, the model’s performance exhibited a gradual decline in the general scenario, while showing an improvement in the hard negative scenario. This led us to hypothesize that an emphasis on hard negatives during training might result in overfitting. Moreover, this finding prompted a reconsideration of how re-ranking models should be evaluated. Relying solely on paired queries for assessment may not suffice, as it risks producing misleading comparisons among different models. This is particularly pertinent because the practical use of the re-ranking model in the end-to-end inference process mirror hard negative scenarios more closely. Our stance is further substantiated by the results of human auditing, detailed in Section 5.5. Here, we observed that evaluations in hard negative contexts correlated more closely with the outcomes of relevance auditing, underscoring the practical applicability of our model in realistic settings.
@@
Index Model NDCG@3 NDCG@3
@@
rr1 Baseline_previous 0.81487 0.30102
rr2 Baseline 0.89459 0.34844
rr3 +ANCE 0.84462 0.64612
rr4 +pairwise loss 0.79172 0.67733
@@ 5.5. Offline query-query relevance auditing
Unfortunately, the quantitative metrics such as NDCG and Recall fail to offer this interpretability as they solely provide empirical insights and lack the depth of an explanatory gold standard. Therefore, this section narrows its focus to query relevance evaluation, emphasizing a human-audited approach over evaluating the model’s compatibility with the testing dataset extracted from historical customer behavior. The auditing protocol designed to assess the semantic relevance between queries within the scope of QR is outlined in Appendix B. In summary, we categorize the query-query relevance into three classes - strict relevant, somewhat relevant, and irrelevant.
@@
index (strictly relevant) (not relevant) correlation
@@
rr2 $0.71178$ $0.57035$ $0.0572$
rr4 $0.77072$ $0.71539$ $0.4254$
From the findings presented in Table 5, a noticeable distinction between the rt1 and rt6 models, as well as between rr2 and rr4 models emerges. The rt1/rr2 model showcased an inability to discern between relevant and irrelevant queries in the hard negative scenario, as indicated by both AUROC and Spearman correlation measures, whereas the rt6/rr4 model exhibited an opposing performance. This outcome suggests that the proposed techniques fostered robustness of QR through bolstering the model’s capacity in accurately modeling query relevance.
In addition, we noticed that the retrieval model rt6 outperformed the re-ranking model rr4 in terms of both AUROC and Spearman correlation metrics. We attribute this disparity to the distinct training objectives of the two models (i.e., JSD for retrieval and KLD for re-ranking). Furthermore, our auditing protocol is more in tune with the requirements of retrieval, which might explain the observed differences in performance. Please refer to Section 4.1 for a detailed exposition.
@@ 5.6. Online experiment
where $\alpha$ and $\beta$ are hyper-parameters. In this application, we conducted two-week A/B tests in different markets, characterized by Japanese, Hindi, and English, respectively, to evaluate the QR implementation introduced in this study. Specifically, for the Japanese case, we conducted two experiments successively with the implementation suggested by the previous work (rt1+rr1) and our proposed implementation (rt6+rr4) as the treatment, respectively, and the system without any QR solution as control; for the Hindi case, we tested the proposed implementation (rt6+rr4) as treatment and the system without QR as control; for the English case, we tested the proposed implementation (rt6+rr4) as treatment while the previous work (rt1+rr1) as control. The results are detailed in Table 6, revealing that the proposed implementation resulted in a considerable improvement in both revenue and conversion in all tested markets. By comparing the previous implementation versus the new implementation in the Japanese (i.e., rt6+rr4 row verses rt1+rr1 row) and English (i.e., rt6+rr4 row) cases, we can further assert that the proposed methods on model training data mining (such as positive query pair labeling and similarity scoring), training objectives (specifically the addition of a sample importance term), and training strategies (such as employing ANCE) are pivotal components for the success of QR practice in terms of scalability and extensibility.
The second application pertains to product search Ads matching. In our experiment performed in Japanese market, we included the Ads matched for $\{q_{i1},...,q_{iN}\}$ directly into the match set of $q_{i}$. The experiment encompassed both QR implementations, one with the proposed techniques (indexed as rt6 and rr4), and another without them (indexed as rt2 and rr2) through separate two-week A/B tests. Table 6 presents the results, indicating that the model with hard negative-tailored adjustments exhibited a noteworthy gain in both Ads revenue and click-through-rate (CTR), while the model without these adjustments showed only a marginal gain. This outcome strongly suggests that effectively handling hard negatives is a crucial component for the successful practice of QR.
@@ 6. Conclusion
Appendix A Statistics of dataset generated in this study
Table 7. Dataset generated in this study.
#unique queries
#positive pairs
#negative pairs
Training+ Validation
5,107,314
339,895,694
N/A
Testing
5,049
371,238
N/A
ANCE (round 1)
5,107,314
339,895,694
743,736,086
ANCE (round 2)
5,107,314
339,895,694
701,336,588
ANCE (round 3)
5,107,314
339,895,694
684,852,381
The statistics of the generated dataset is shown in Table 7. It is worth mentioning that $(q_{i},q_{j})$ and $(q_{j},q_{i})$ are counted as two separate pairs in the counting. We also provide the information on the ANCE dataset. Notabally, the ANCE dataset demonstrates a trend toward convergence across iterations. In the second round, it retains 33.8% of hard negative query pairs from the first round, while in the third round, it retains 64.6% of the pairs from the second round. We only perform three rounds of ANCE as the empirical result shows marginal benefit with more rounds.
Appendix B Human audit protocol
The human auditor is tasked with categorizing each query-query pair into one of three following classes.
Strict relevant - Query 1 and query 2 are different representations of the same shopping intent (i.e., customer would expect the search result of query 1 and query 2 to largely overlap with each other).
Somewhat relevant - The shopping intent of query 1 and query 2 have no conflict but are not exactly the same, for example, query 1’s shopping intent is much broader and entirely cover query 2’s, or the opposite case (i.e., customer expect to find item related to query 2 when search for query 1, or vice versa. But the overlap between search result of query 1 and 2 is only a small portion of the search result of query 1 or 2).
Not relevant - Query 1 and query 2 have some related components, but share conflict shopping intent in either of 1) Product Type, 2) Model, 3) Age/gender group, 4) Size/color.

### v2

@@ 4.5. Model training with hard negative mining
Previous research has introduced the concept of Approximate Nearest Neighbor Negative Contrastive Estimation (ANCE) (Sohn, 2016) to address the inefficacy of using in-batch negative in contrastive learning. ANCE entails the incorporation of hard negatives, which are selected from the entire dataset for each query within a batch. This selection process is carried out through a sequence of retrieval model inferences and KNN indexing. In the context of QR, we define the set of hard negatives as the queries that are retrieved for $q_{i}$ but do not share any co-purchases with $q_{i}$.
@@
We put forth distinct strategies to harness ANCE in the training of both retrieval and re-ranking models. For the retrieval model, we employ a self-learning strategy, wherein the bi-encoder is iteratively finetuned with the ANCE dataset it generates. Conversely, for the re-ranking model, as it is hard to acquire hard-negatives by cross-encoder itself through a retrieval process, we adopt a learning-from-teacher strategy. Under this approach, the cross-encoder is finetuned using the ANCE dataset generated by the bi-encoder model. We will highlight in the experiment section that this approach is necessary regarding re-ranking model’s performance. In addition, we group the pairs associated to each $q_{i}$ to form a positive set and employ pairwise loss rather than pointwise loss. The necessity for training in a pairwise manner arises particularly in the case of hard negatives, where positive and negative samples often exhibit only nuanced differences in their text representations. Pairwise training can enable the model to capture the supervised signal with greater precision, reduce variance during training, and thus, enhance the overall training efficacy. Specifically, we use circle loss (Sun et al., 2020) as training objective.
@@
where $s_{CE}(q_{i},q_{j})$ represents the similarity score given by the cross-encoder model, and $H(q_{i})$ represents the hard negatives of $q_{i}$. The ANCE mining and model continuously fine-tuning process is described in Figure 3.
Figure 3. The ANCE mining and model continuous finetune process. Specifically, the bi-encoder model is iteratively finetuned using ANCE data derived by self-retrieved pairs subtracting pairs with co-purchase. Then, the cross-encoder model is finetuned with the ANCE data from the bi-encoder.
@@ 5.1. Dataset generation
We construct training and testing dataset with one-year anonymized Amazon search data. We first extract query-product pairs from anonymized historical customer behavior data. Then we perform self-join of these query-product pairs on product, and aggregate the joined result by product to obtain the query co-purchase data. At last, we select 5,049 unique queries from $D$ and extract all query pairs containing the selected queries as testing data (abour 371K). The remaining pairs are divided into training and validation data, using a 90%/10% ratio.
@@ 5.2. Offline evaluation criteria
We employ distinct evaluation metrics for the retrieval and re-ranking tasks. For retrieval, we calculate the recall@100 for both the top-3 paired queries and all queries paired to the given query in the testing data. This separation into top-3 and all queries is motivated by their significance in reflecting the model’s performance in downstream ranking and matching applications, respectively. Besides, we calculate both micro-recall, which represents the average recall of each query in the testing dataset, and macro-recall, which is derived by dividing the total number of retrieved paired queries by the total number of all paired queries.
@@ 5.3. Offline experiment result
We train a series of bi-encoder model and cross-encoder model, separately for retrieval and re-ranking tasks. The proposed approaches are progressively added upon a baseline model to study the individual benefit brought by each approach. The datasets used for training were constructed through query normalization-based augmentation.
Retrieval - The experimental results are detailed in Table 3. The baseline model ( rt1) did not employ sample weighting; rather, it used the top 30% of pairs for each query as training data. For simplicity, we will analyze the findings specifically for recall@100 in relation to the top-3 queries (in bold font), as the results for all metrics follow a similar trend. Notably, the inclusion of the sample importance term elevated recall@100 by 5.4% ( rt2 vs. rt1). This improvement was further augmented by the addition of ANCE, resulting in an incremental gain of 7.9% ( rt3 vs. rt2). Additionally, through successive iterations of the ANCE dataset and continuous model fine-tuning, recall@100 increased to 0.777 in round 2 ( rt4) and 0.798 in round 3 ( rt5). Note that we only perform three rounds of ANCE as we observed marginal benefit with more rounds. The outcomes suggest that the combined utilization of the sample importance in the loss function and ANCE substantially enhances the retrieval model’s performance.
@@
Model Recall (top-3) Recall
@100 @100
micro macro micro macro
rt1 Baseline 0.5903 0.5932 0.4701 0.3451
rt2 +sample importance 0.6441 0.9322 0.5118 0.3788
rt3 +ANCE (round 1) 0.7232 0.7256 0.6075 0.4638
rt4 +ANCE (round 2) 0.7766 0.7771 0.6570 0.4876
rt5 +ANCE (round 3) 0.7984 0.7995 0.6808 0.5159
Re-ranking - the model performance is presented in Table 4. The baseline model trained with the KLD objective outperformed the best retrieval model (rt5), indicating the necessity of the re-ranking stage under the QR context. Surprisingly, our findings further revealed an unexpected outcome: as the model’s focus shifted towards hard negative samples through the incorporation of ANCE and the adoption of circle loss, the model’s performance exhibited a gradual decline in the general scenario, while showing an improvement in the hard negative scenario. This led us to hypothesize that an emphasis on hard negatives during training might result in overfitting. Moreover, this finding prompted a reconsideration of how re-ranking models should be evaluated. Relying solely on paired queries for assessment may not suffice, as it risks producing misleading comparisons among different models. This is particularly pertinent because the practical use of the re-ranking model in the end-to-end inference process mirror hard negative scenarios more closely. Our stance is further substantiated by the results of human auditing, detailed in Section 5.4. Here, we observed that evaluations in hard negative contexts correlated more closely with the outcomes of relevance auditing, underscoring the practical applicability of our model in realistic settings.
@@
Model NDCG@3 NDCG@3
@@
rt5 Retrieval 0.81487 0.30102
rr1 Baseline 0.89459 0.34844
rr2 +ANCE 0.84462 0.64612
rr3 +pairwise loss 0.79172 0.67733
@@ 5.4. Offline query-query relevance auditing
(strictly relevant) (not relevant) correlation
@@
rr1 $0.71178$ $0.57035$ $0.0572$
rr3 $0.77072$ $0.71539$ $0.4254$
From the findings presented in Table 5, a noticeable distinction between the rt1 and rt5 models, as well as between rr1 and rr3 models emerges. The rt1/rr1 model showcased an inability to discern between relevant and irrelevant queries in the hard negative scenario, as indicated by both AUROC and Spearman correlation measures, whereas the rt5/rr3 model exhibited an opposing performance. This outcome suggests that the proposed techniques fostered robustness of QR through bolstering the model’s capacity in accurately modeling query relevance.
In addition, we noticed that the retrieval model rt6 outperformed the re-ranking model rr3 in terms of both AUROC and Spearman correlation metrics. We attribute this disparity to the distinct training objectives of the two models (i.e., JSD for retrieval and KLD for re-ranking). Furthermore, our auditing protocol is more in tune with the requirements of retrieval, which might explain the observed differences in performance. Please refer to Section 4.3 for a detailed exposition.
@@ 5.5. Online experiment
where $\alpha$ and $\beta$ are hyper-parameters. In this application, we conducted two-week A/B tests in different markets, characterized by Japanese, Hindi, and English, respectively, to evaluate the QR implementation introduced in this study. The results are detailed in Table 6, revealing that the proposed implementation resulted in a considerable improvement in both revenue and conversion in all tested markets. By comparing the previous implementation versus the new implementation in the Japanese, Hindi (i.e., rt5+rr3 row verses rt1+rr1 row), and English (i.e., rt5+rr3 row) cases, we can further assert that the proposed refinements to model training are pivotal components for the success of QR practice in terms of scalability and extensibility.
The second application pertains to product search Ads matching. In our experiment performed in Japanese market, we included the Ads matched for $\{q_{i1},...,q_{iN}\}$ directly into the match set of $q_{i}$. The experiment encompassed both QR implementations, one with the proposed techniques (rt5+rr3), and another without them (rt1+rr1) through separate two-week A/B tests. Table 6 presents the results, indicating that the model with hard negative-tailored adjustments exhibited a noteworthy gain in both Ads revenue and click-through-rate (CTR), while the model without these adjustments showed only a marginal gain. This outcome strongly suggests that effectively handling hard negatives is a crucial component for the successful practice of QR.